### v1

# HalluGuard: Demystifying Data-Driven and Reasoning-Driven Hallucinations in LLMs

Abstract
The reliability of Large Language Models (LLMs) in high-stakes domains such as healthcare, law, and scientific discovery is often compromised by hallucinations. These failures typically stem from two sources: data-driven hallucinations and reasoning-driven hallucinations. However, existing detection methods usually address only one source and rely on task-specific heuristics, limiting their generalization to complex scenarios. To overcome these limitations, we introduce the Hallucination Risk Bound, a unified theoretical framework that formally decomposes hallucination risk into data-driven and reasoning-driven components, linked respectively to training-time mismatches and inference-time instabilities. This provides a principled foundation for analyzing how hallucinations emerge and evolve. Building on this foundation, we introduce HalluGuard, a NTK-based score that leverages the induced geometry and captured representations of the NTK to jointly identify data-driven and reasoning-driven hallucinations. We evaluate HalluGuard on 10 diverse benchmarks, 11 competitive baselines, and 9 popular LLM backbones, consistently achieving state-of-the-art performance in detecting diverse forms of LLM hallucinations.
1 Introduction
Large language models (LLMs) are increasingly deployed in high-stakes domains such as healthcare, law, and scientific discovery(Bommasani et al., 2021; Thirunavukarasu et al., 2023). However, adoption in these settings remains cautious, as such domains are highly regulated and demand strict compliance, interpretability, and safety guarantees(Dennstädt et al., 2025; Kattnig et al., 2024). A major barrier is the risk of hallucinations, generated content appears unfaithful or nonsensical. Such errors can have severe consequences(Dennstädt et al., 2025)—as the example in Figure 1, a generated incorrect medical diagnosis may delay treatment or lead to harmful interventions. Therefore, detecting hallucinations is not merely a technical challenge but a prerequisite for trustworthy deployment, as undetected errors undermine reliability, accountability, and user safety.
Generally, hallucinations in LLMs arise from two primary sources(Ji et al., 2023; Huang et al., 2023): data-driven hallucinations, which stem from flawed, biased, or incomplete knowledge encoded during pre-training or fine-tuning; and reasoning-driven hallucinations, which originate from inference-time failures such as logical inconsistencies or breakdowns in multi-step reasoning(Zhang et al., 2023; Zhong et al., 2024). Detection methods broadly split along these two dimensions. Approaches for data-driven hallucinations often compare outputs against retrieved documents or references(Shuster et al., 2021; Min et al., 2023; Ji et al., 2023), or exploit sampling consistency as in SelfCheckGPT(Manakul et al., 2023). In contrast, methods for reasoning-driven hallucinations rely on signals of inference-time instability, including probabilistic measures such as perplexity(Ren et al., 2022), length-normalized entropy(Malinin and Gales, 2020), semantic entropy(Kuhn et al., 2023), energy-based scoring(Liu et al., 2020), and RACE(Wang et al., 2025). Others probe internal representations, for example, Inside(Chen et al., 2024a), which applies eigenvalue-based covariance metrics and feature clipping, ICR Probe(Zhang et al., 2025), which tracks residual-stream updates, and Shadows in the Attention(Wei et al., 2025), which analyzes representation drift under contextual perturbations. While these methods shed light on the mechanisms underlying hallucinations, most remain tailored to a single hallucination type and fail to capture their evolution. Yet growing evidence indicates that data-driven and reasoning-driven hallucinations often evolve during multi-step generation(Liu et al., 2025; Sun et al., 2025). As shown in Figure 1, it emerges from an initial disease misclassification and evolves into a distorted diagnosis, delaying treatments and risking fatality. This gap brings two central questions: (1) How can we develop a unified theoretical understanding of how hallucinations evolve? and (2) How can we detect them effectively and efficiently without relying on external references or task-specific heuristics?
To address these challenges, we propose a unified theoretical framework–Hallucination Risk Bound, which decomposes the overall hallucination risk into two components: a data-driven term, capturing semantic deviations rooted in inaccurate, imbalanced, or noisy supervision acquired during model training; and a reasoning-driven term, reflecting instability introduced by inference-time dynamics, such as logical missteps or temporal inconsistency. This decomposition not only elucidates the mechanism behind hallucinations but also reveals how they emerge and evolve. Specifically, our analysis shows that hallucinations originate from semantic approximation gaps-captured by representational limits of the model-and are subsequently amplified by unstable rollout dynamics, evolving across decoding steps. As such, our framework offers a unified theoretical lens for characterizing the emergence and evolution of these hallucinations.
Figure 1: An illustration of hallucination emerging and evolving in the context of disease diagnosis.
Building on the theoretical foundation, we propose HalluGuard, a Neural Tangent Kernel(NTK)-based score that leverages the induced geometry and captured representations of the NTK to jointly identify data-driven and reasoning-driven hallucinations. We evaluate HalluGuard comprehensively across 10 diverse benchmarks, 11 competitive baselines, and 9 popular LLM backbones. HalluGuard consistently achieves state-of-the-art hallucination detection performance, demonstrating its efficacy.
2 Preliminaries
Hallucination Detection.
There are two primary sources of hallucinations in LLMs(Ji et al., 2023; Huang et al., 2023): data-driven hallucination, which stems from incomplete or biased knowledge encoded during pre-training or fine-tuning, and reasoning-driven hallucination, which arises from unstable or inconsistent inference dynamics at decoding time. This distinction has implicitly guided a broad range of detection strategies, which we examine through these two lenses.
For data-driven causes, a recurring signal is elevated predictive uncertainty. A common formulation adopts the sequence-level negative log-likelihood:
$\mathcal{U}(\mathbf{y}\mid\mathbf{x},\theta)=-\frac{1}{T}\sum_{t=1}^{T}\log p_{\theta}(y_{t}\mid y_{<t},\mathbf{x}),$ (1)
which quantifies the average uncertainty of generating a sequence $\mathbf{y}=[y_{1},\ldots,y_{T}]$ from input $\mathbf{x}$ and $\theta$ denotes model parameters. This directly recovers Perplexity(Ren et al., 2022), where low scores imply confident predictions, while high scores indicate implausible generations due to weak priors. To capture more nuanced uncertainty, later methods extend this formulation to multi-sample settings. The Length-Normalized Entropy(Malinin and Gales, 2020) penalizes dispersion across stochastic generations $\mathcal{Y}=\{\mathbf{y}^{1},\ldots,\mathbf{y}^{K}\}$, offering a finer-grained view of model indecision. This perspective is further enriched by Semantic Entropy(Kuhn et al., 2023), which projects sampled responses into semantic space, and by energy-based scoring(Liu et al., 2020), which replaces log-probability with a learned confidence function. Collectively, these methods reflect a progression from token-level likelihoods to semantically grounded multi-sample uncertainty estimators.
In contrast, reasoning-driven hallucinations arise from brittle inference trajectories, where identical contexts may yield inconsistent or incoherent outputs. A commonly used measure of such instability is the cross-sample consistency score:
$\mathcal{C}(\mathcal{Y}\mid\mathbf{x},\theta)=\frac{1}{C}\sum_{i=1}^{K}\sum_{j=i+1}^{K}\text{sim}(\mathbf{y}^{i},\mathbf{y}^{j}),\vskip-5.0pt$ (2)
where $C=K\cdot(K-1)/2$, and $\text{sim}(\cdot,\cdot)$ is a similarity function such as ROUGE-L(Lin, 2004), cosine similarity, or BLEU(Chen et al., 2024b). Low scores reflect diverging generations and unstable reasoning. Several reasoning-driven detection methods can be interpreted through this lens. Early approaches used surface-level lexical overlap metrics(Lin et al., 2022b), while SelfCheckGPT(Manakul et al., 2023) advanced this by evaluating factual entailment across responses, and FActScore(Min et al., 2023) extended this further by comparing outputs to retrieved reference documents. More recent efforts probe internal signals directly: Inside(Chen et al., 2024a) analyzes the covariance spectrum of embedding representations, and RACE(Wang et al., 2025) diagnoses instability in multi-step reasoning.
NTK in LLMs.
NTK provides a principled framework for analyzing the training dynamics in the overparameterized regime characteristic of modern LLMs(Jacot et al., 2020). Formally, for a network output $f(x,\theta)$ with input $x$ and parameters $\theta$, the NTK is defined as:
$\Theta(x,x^{\prime},\theta)=\nabla_{\theta}f(x,\theta)\cdot\nabla_{\theta}f(x^{\prime},\theta).$ (3)
This kernel $\Theta(x,x^{\prime},\theta)$ quantifies the similarity of training dynamics between inputs $x$ and $x^{\prime}$. In the infinite-width limit, it converges to a deterministic value at initialization and remains nearly constant throughout training(Lee et al., 2020b). This stability reduces the highly nonlinear optimization of deep networks to a tractable kernel regression problem. By examining the eigenspectrum of the NTK, one can probe how internal representations are shaped during training: which features are prioritized (e.g., syntax versus semantics), how quickly different tasks converge, and why overparameterized networks generalize effectively to unseen data(Ju et al., 2022). In this way, the NTK transforms the apparent complexity of LLM optimization into a clear lens on how these models capture, process, and generalize information(Zeng et al., 2025).
3 Methodology
3.1 Problem Setting
Our analysis reveals that hallucination is not a unified failure mode but rather shifts with the task structure. On the instruction-following Natural benchmark(Wang et al., 2022), 88.9% of the overall 3499 errors are from logical missteps (reasoning-driven) while 11.1% are factual inaccuracies (data-driven). By contrast, on the math-focused MATH-500(Hendrycks et al., 2021), the 1985 wrong generations are dominated by 1946 reasoning errors (98.1%), with only 19 factual flaws (1.9%). This contrast highlights that, in practice, hallucinations are rarely pure but often mixtures of data-driven bias and reasoning-driven instability—motivating our formal decomposition of hallucination sources.
Problem Definition. Let $\mathcal{Y}$ denote the space of textual outputs and let $\Phi:\mathcal{Y}\rightarrow U_{h}$ be a task-specific encoder that maps textual sequences into the hypothesis space $U_{h}$, equipped with a norm $\|\cdot\|$ (e.g., task-calibrated embedding space or structured metric). We interpret each $u\in U_{h}$ as a reasoning chain, composed of step-wise logical statements. For an input $\mathbf{x}$ with ground-truth output $y^{*}\in\mathcal{Y}$, define the gold-standard reasoning chain as $u^{*}:=\Phi(y^{*})\in U_{h}$. An LLM with parameters $\theta$ emits a random sequence $Y=(Y_{1},\dots,Y_{T})$ via $p_{\theta}(y_{t}\mid y_{<t},\mathbf{x})$, yielding a predicted reasoning chain $u_{h}:=\Phi(Y)\in U_{h}$. Its expected value under the model’s decoding distribution is $\mathbb{E}[u_{h}]:=\mathbb{E}_{Y\sim p_{\theta}(\cdot\mid\mathbf{x})}[\Phi(Y)].$
We consider perturbations in a local neighborhood of the decoding process. Let $\delta\in\mathbb{R}^{r}$ parameterize a small perturbation (e.g., of the prefix tokens, step-$t$ logits, or hidden state), and let $\mathcal{B}_{\rho}:=\{\delta:\|\delta\|\leq\rho\}$. Define the perturbed decoder map $G:\mathbb{R}^{r}\!\to\!U_{h}$ by $G(\delta):=\Phi\!\big(Y(\delta)\big)$, where $Y(\delta)$ is the sequence under perturbation. Let $J\in\mathbb{R}^{d_{h}\times r}$ denote the (Gauss–Newton) Jacobian of $G$ at $\delta=0$. Our goal is to formalize how hallucination emerges and evolves in LLMs.
3.2 Hallucination Risk Bound
To bridge the formal setup with the phenomenon of hallucination, we first disentangle the sources of hallucinations. Intuitively, hallucinations may arise either from systematic biases in the knowledge encoded by the model (data-driven) or from instabilities during autoregressive decoding (reasoning-driven). The following proposition formalizes this idea by decomposing the total hallucination risk into two components.
We first impose the following assumptions:
  • A1.
    $(U,\|\cdot\|)$ is a Hilbert space; $\Phi$ is measurable with unique best solution and $\|\Phi(Y)\|$ has finite second moment.
  • A2.
    Triangle inequality holds for $\|\cdot\|$ and $\Phi$ is $L_{\Phi}$-Lipschitz w.r.t. an edit distance on $\mathcal{Y}$.
  • A3.
    For $\delta\in\mathcal{B}_{\rho}$, the mapping $G$ admits the local expansion $G(\delta)=G(0)+J\delta+R(\delta)$, where the remainder is bounded by $\|R(\delta)\|\leq\tfrac{1}{2}H_{\star}\|\delta\|^{2}$ for some curvature constant $H_{\star}>0$.
Proposition 3.1 (Hallucination Risk Decomposition). Under A1–A3, applying the triangle inequality yields a natural split of the risk: $\|u^{*}-u_{h}\|\;\leq\;\underbrace{\|u^{*}-\mathbb{E}[u_{h}]\|}_{\text{data-driven term}}\;+\;\underbrace{\|u_{h}-\mathbb{E}[u_{h}]\|}_{\text{reasoning-driven term}}$ This decomposition distinguishes errors caused by systematic bias in the learned representation from those introduced during stochastic rollout.
Characterizing Data-Driven Hallucination.
To quantify the data-driven term, we take inspiration from the NTK, which has proven effective in analyzing training dynamics of overparameterized models. Here, NTK geometry provides a way to measure how well the model’s representation space aligns with task generation under small perturbations.
Let $U_{h}\subset U$ denote the hypothesis subspace accessible to the model under perturbations. By Céa’s lemma(Céa, 1964) with curvature penalty, the data-driven term can be bounded as
$\|u^{*}-\mathbb{E}[u_{h}]\|\;\leq\;\frac{\Lambda}{\gamma}\inf_{u\in U_{h}}\|u^{*}-u\|,$ (4)
where $\gamma=\lambda_{\min}(\mathcal{K}_{\Phi})$ is the smallest eigenvalue of the NTK Gram matrix on embedded perturbations, and $\Lambda\leq\|T\|$ reflects the operator norm of the problem/operator mapping $\mathcal{T}$. Intuitively, the ratio $\tfrac{\Lambda}{\gamma}$ measures the conditioning of the feature map: well-conditioned NTK spectra allow a closer approximation to the true generation.
This ratio can be further controlled in terms of pretraining–finetuning mismatch:
$\frac{\Lambda}{\gamma}\;\leq\;1\;+\;k_{\mathrm{pt}}\log\!\mathcal{O}(P,L)\;+\;k\cdot\frac{\epsilon_{\mathrm{mismatch}}}{\mathrm{Signal}_{k}},$ (5)
where $\log\!\mathcal{O}(P,L)$ is a complexity term from parameter count $P$ and prompt length $L$, $\epsilon_{\mathrm{mismatch}}$ denotes the Wasserstein distance between prompt and query distributions, $\mathrm{Signal}_{k}$ measures task-aligned energy in the top-$k$ eigenspace. $k_{\mathrm{pt}}$ and $k$ are task and model-dependent constants. Thus, data-driven hallucinations grow when the mismatch is large or when the task signal is weak.
Characterizing Reasoning-Driven Hallucination.
The reasoning-driven term captures reasoning-driven instability that accumulates during autoregressive decoding. Here, we model generation as a martingale process, where deviation from the expectation is controlled by concentration inequalities. Specifically, Freedman’s inequality(Geman et al., 1992) gives
$\|u_{h}-\mathbb{E}[u_{h}]\|\;\leq\;K\cdot\exp\!\Big(-\tfrac{K\epsilon^{2}}{C}\Big)\cdot\alpha(e^{\beta T}-1),$ (6)
where $K$ is the number of rollouts averaged, $\beta$ summarizes per-step growth in local Jacobians, $\alpha$ scales the cumulative effect and $C$ is a task and model-dependent constant. This bound shows that reasoning-driven hallucinations grow exponentially with sequence length $T$.
We now synthesize the two components into a unified result that characterizes the overall risk of hallucination. By combining the NTK-conditioned approximation bound for data-driven deviation with the Freedman-style concentration bound for reasoning-driven instability, we obtain the following unified bound of data-driven and reasoning-driven hallucinations (detailed proof is provided in Appendix A):
Theorem 3.2 (Hallucination Risk Bound). Let $u^{*}:=\Phi(y^{*})$ denote the semantic embedding of the ground-truth output and $u_{h}:=\Phi(Y)$ that of the model-generated output. Under Assumptions A1–A3, suppose there exists $\beta\geq 0$ such that $\Big\|\prod_{t=1}^{T}J_{t}\Big\|_{2}\;\leq\;e^{\beta T}.$ Then the total hallucination risk satisfies $\|u^{*}-u_{h}\|\;\leq\;\underbrace{\Big(1+k_{\mathrm{pt}}\log\mathcal{O}(P,L)+k\cdot\tfrac{\epsilon_{\mathrm{mismatch}}}{\mathrm{Signal}_{k}}\Big)\inf_{u\in U_{h}}\|u^{*}-u\|}_{\text{data-driven term}}\;+\;\underbrace{|\mathcal{L}|\cdot\exp\!\Big(-\tfrac{K\epsilon^{2}}{C}\Big)\cdot\alpha\big(e^{\beta T}-1\big)}_{\text{reasoning-driven term}}$
3.3 Hallucination Quantification via HalluGuard
While Theorem 3.2 makes explicit how data-driven and reasoning-driven hallucinations emerge and evolve, applying it directly at inference is impractical since direct step-wise Jacobians for billion-parameter LLMs are intractable, so we seek a proxy score that is computable, stable, and faithful to our decomposition.
Let $\mathcal{K}$ denote the NTK Gram matrix with eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{r}>0$ and condition number $\kappa(\mathcal{K})=\lambda_{\max}/\lambda_{\min}$. Let $J_{t}$ be the step-$t$ input–output Jacobian of the decoder, and define $\sigma_{\max}:=\sup_{t}\|J_{t}\|_{2}$ as the uniform spectral bound(note that $\sigma_{\max}$ is independent of the spectrum of $\mathcal{K}$).
Under Assumptions A1–A3, a standard NTK approximation argument yields $\inf_{u\in U_{h}}\|u^{*}-u\|\;\leq\;C_{d}\,\det(\mathcal{K})^{-c_{d}}\,\|u^{*}\|,$ so that $\det(\mathcal{K})$ capture the representations in systematic bias.
For autoregressive rollout, based on the property of Jacobian, we have $\Big\|\prod_{t=1}^{T}J_{t}\Big\|_{2}\;\leq\;\prod_{t=1}^{T}\|J_{t}\|_{2}\;=\;\exp\!\Big(\sum_{t=1}^{T}\log\|J_{t}\|_{2}\Big)$ , so that we have $\Big\|\prod_{t=1}^{T}J_{t}\Big\|_{2}\;\leq\;e^{\beta T}.$ Since $\beta\leq\log\sigma_{\max}$ with $\sigma_{\max}:=\sup_{t}\|J_{t}\|_{2}$ thus we have the upper bound as $\|\prod_{t=1}^{T}J_{t}\|_{2}\leq\sigma_{\max}^{T}=e^{(\log\sigma_{\max})T}$. Thus, $\log\sigma_{\max}$ serves as a stable and tractable proxy for the per-step amplification rate.
Perturbation analysis of $\mathcal{K}$, together with classical eigenvalue sensitivity results(Trefethen and Bau, 2022), yields $\mathrm{Var}[u_{h}]\;\leq\;c_{v}\,\kappa(\mathcal{K})^{2}\,\|\delta\|^{2},$ showing that instability grows quadratically with the condition number $\kappa(\mathcal{K})$. To temper this effect and ensure additivity, we penalize ill-conditioned representations via $\log\kappa^{2}$, where log compression brings a well-behaved dynamic range.
Table 1: Correlation between NTK proxies and task families.
SQuAD Math-500 TruthfulQA
$\det(\mathcal{K})$ 0.84 0.42 0.61
$\log\sigma_{\max}-\log\kappa^{2}$ 0.39 0.88 0.67
In summary, $\det(\mathcal{K})$ quantifies representational adequacy, $\log\sigma_{\max}$ captures rollout amplification, and $\log\kappa^{2}$ penalizes spectral instability, together forming a compact and tractable proxy consistent with the Hallucination Risk Bound. The lightweight projection layers are self-supervised spectral calibration modules, optimized offline (via AdamW) to align NTK spectral properties across heterogeneous backbones into a stable, comparable geometric space—without hallucination labels or task-specific supervision, with the backbone fully frozen and zero runtime overhead during inference. Detailed proofs are provided in Appendix B.
Empirical validation.
We empirically validate how those proxies correlate with different task families. In Table 1, $\det(\mathcal{K})$ correlates most strongly with the data-centric task SQuAD ($0.84$), indicating its role in capturing factual fidelity. In contrast, for the reasoning-oriented MATH-500, the highest correlation is observed with $\log\sigma_{\max}-\log\kappa^{2}$ ($0.88$), reflecting the importance of amplification and stability in multi-step reasoning.
Motivated by the above, we formally define HalluGuard as follows, which provides a principled and unified lens for hallucination detection:
$\boxed{\text{{{HalluGuard}}}(u_{h})\;=\;\det(\mathcal{K})\;+\;\log\sigma_{\max}\;-\;\log\kappa^{2}.}$ (7)
4 Experiments
We comprehensively evaluate HalluGuard across 10 diverse benchmarks, 11 competitive baselines, and 9 popular LLM backbones. We aim to evaluate its efficacy from the following five questions: Q1: How does HalluGuard perform across different task families? Q2: How does HalluGuard perform across LLMs of different scales? Q3: How does each term capture trends across task families? Q4: Can HalluGuard guide test-time inference to improve downstream reasoning? Q5: How well does HalluGuard generalize to detecting fine-grained hallucinations beyond benchmarks?
Section 4.1 details the setup; Section 4.2 evaluates HalluGuard as a detection method(Q1–Q3), Section 4.3 applies HalluGuard in score-guided inference(Q4) and Section 4.4 analyzes HalluGuard on fine-grained hallucination via a case study on semantic data(Q5).
4.1 Evaluation Setup
Benchmarks. We evaluate across 10 widely used benchmarks spanning three distinct categories. For data-grounded QA, we include RAGTruth(Niu et al., 2024), NQ-Open(Kwiatkowski et al., 2019), HotpotQA(Yang et al., 2018) and SQuAD(Rajpurkar et al., 2016), which emphasize factual correctness through external evidence. For reasoning-oriented tasks, we use GSM8K(Cobbe et al., 2021), MATH-500(Hendrycks et al., 2021), and BBH(Suzgun et al., 2022), which require multi-step derivations prone to compounding errors. Finally, for instruction-following settings, we consider TruthfulQA(Lin et al., 2022a), HaluEval(Li et al., 2023) and Natural(Wang et al., 2022), which probe hallucinations under open-ended or adversarial prompts.
Baselines. We compare HalluGuard with 11 competitive detectors spanning diverse strategies. Uncertainty-based methods include Perplexity(Ren et al., 2022), Length-Normalized Predictive Entropy(LN-Entropy)(Malinin and Gales, 2020), Semantic Entropy(Kuhn et al., 2023), Energy Score(Liu et al., 2020) and P(true)(Kadavath et al., 2022). Consistency-based approaches cover SelfCheckGPT(Manakul et al., 2023), Lexical Similarity(Lin et al., 2022b), FActScore(Min et al., 2023) and RACE(Wang et al., 2025). Internal-state methods are represented by Inside(Chen et al., 2024a) and MIND(Su et al., 2024).
LLM Backbone Models. We evaluate 9 publicly available LLMs spanning different scales and architectures. These include five models from the Llama family (Llama2-7B, Llama2-13B, Llama2-70B, Llama3-8B, and Llama3.2-3B)(Touvron et al., 2023; Grattafiori et al., 2024), along with OPT-6.7B(Zhang et al., 2022), Mistral-7B-Instruct(Jiang et al., 2023), QwQ-32B(Yang et al., 2024), and GPT-2 (117M)(Radford et al., 2019). All models are used in their off-the-shelf form with pre-trained weights and tokenizers provided by Hugging Face, without further fine-tuning.
Evaluation Metrics. We evaluate hallucination detection ability under two regimes following Janiak et al. (2025): ROUGE-based reference evaluation ($_{r}$) and LLM-as-a-Judge ($_{\mathrm{llm}}$). For performance measures, we report the area under the receiver operating characteristic curve (AUROC) and the area under the precision–recall curve (AUPRC). AUROC is widely used to assess the quality of binary classifiers and uncertainty estimators, while AUPRC highlights performance under class imbalance. In both cases, higher values indicate better detection.
4.2 Main Results
Q1: How does HalluGuard perform across different task families? To evaluate how HalluGuard performs across different task types, we conduct experiments on all benchmarks. For clarity, Table 2 presents representative results from three task families: data-centric (RAGTruth), reasoning-oriented (Math-500), and instruction-following (TruthfulQA). As shown, HalluGuard consistently outperforms all baselines across backbones. On Math-500, it reaches 81.76% AUROC and 79.76% AUPRC, improving over the second-best method by up to 8.3%. On RAGTruth, it attains 84.59% AUROC and 81.15% AUPRC, with gains of up to 7.7%. On TruthfulQA, it achieves 77.05% AUROC and 73.79% AUPRC, exceeding the next strongest baseline by as much as 6.2%. Overall, HalluGuard establishes new state-of-the-art results across diverse task families, with particularly pronounced improvements on reasoning-oriented benchmarks.
Table 2: Performance comparison on representative benchmarks: data-centric (RAGTruth), reasoning-oriented (Math-500), and instruction-following (TruthfulQA). We highlight the first and second best results.
GPT2 OPT-6.7B Mistral-7B QwQ-32B
AUROCr
AUPRCr
[…]
AUROCr
AUPRCr
[…]
AUROCr
AUPRCr
[…]
AUROCr
AUPRCr
[…]
RAGTruth HalluGuard 75.51 73.40 62.40 56.60 80.13 76.77 71.01 63.58 82.31 80.79 64.89 67.25 84.59 81.15 71.82 66.68
Inside 73.42 73.08 61.99 56.39 79.49 71.82 66.1 62.46 75.32 73.19 64.58 61.05 77.72 73.47 66.05 64.73
MIND 58.54 54.79 43.47 41.85 63.82 62.58 51.03 44.78 73.13 71.53 58.25 58.6 64.23 63.06 47.37 51.47
Perplexity 58.07 56.68 43.84 41.53 64.47 61.57 47.12 52.98 65.42 63.63 53.28 51.36 73.91 72.92 60.81 59.77
LN-Entropy 64.42 60.79 49.41 45.04 60.81 57.91 48.76 42.27 64.22 60.92 52.24 48.41 63.81 62.26 47.52 52.17
Energy 65.53 62.42 51.8 47.22 66.54 63.28 54.21 49.19 64.36 62.26 48.64 53.93 73.26 71.21 65.43 62.32
Semantic Ent. 60.72 59.41 50.55 45.86 70.2 68.34 54.54 56.74 66.01 64.49 53.01 55.5 66.48 64.41 51.54 50.11
Lexical Sim. 64.72 63.1 55.04 48.04 67.28 64.62 52.55 54.86 64.96 61.17 52.34 45.11 70.87 67.41 61.25 51.01
SelfCheckGPT 65.4 62.79 52.85 52.43 66.64 64.89 52.69 51.17 71.19 68.45 63.13 60.23 65.79 62.45 54.76 51.29
RACE 64.83 62.84 51.8 48.44 64.26 61.03 52.74 46.22 66.34 64.54 51.88 53.86 71.13 69.96 57.58 55.54
P(true) 66.19 64.04 48.2 56.27 68.44 65.48 57.53 53.08 72.54 71.8 57.25 59.42 65.32 63.01 53.01 52.32
FActScore 65.72 64.39 51.94 47.51 61.53 58.2 51.86 45.57 63.98 60.71 53.54 49.34 66.72 64.03 58.21 49.17
BBH HalluGuard 71.06 67.94 62.05 59.05 73.1 70.88 63.67 61.88 79.85 76.5 67.13 60.57 81.76 79.76 68.77 65.46
Inside 66.18 66.81 56.15 58.62 70.64 65.22 63.28 59.28 67.2 65.49 51.3 53.46 80.8 71.49 64.05 63.42
MIND 55.41 51.77 39.01 41.59 55.48 53.46 38.59 40.88 65.71 63.7 49.61 52.54 61.75 60.18 53.46 50.04
Perplexity 53.28 50.22 43.86 38.98 64.89 62.12 48.65 51.99 61.97 60.05 51.15 42.87 60.28 57.75 51.62 43.38
LN-Entropy 60.84 58.76 42.76 47.48 58.71 55.01 43.55 42.02 68.96 69.44 58.79 57.49 63.96 62.18 46.01 49.5
Energy 55.09 51.99 46.2 39.5 53.96 50.98 42.56 34.12 66.27 62.72 49.48 50.06 69.61 68.66 54.35 57.36
Semantic Ent. 58.16 54.81 49.61 40.39 62.63 59.52 50.14 45.02 64.99 61.33 50.11 45.53 62.76 60.95 45.77 45.75
Lexical Sim. 51.37 47.18 38.37 39.06 61.27 58.06 44.13 42.96 58.25 55.92 46.31 46.01 69.46 67.59 55.93 52.6
SelfCheckGPT 54.51 51.86 44.62 44.01 57.36 53.21 42.55 38.27 63.68 62.5 51.7 53.03 64.56 62.49 55.85 45.8
RACE 55.99 54.66 41.39 38.32 64.23 62.03 56.03 53.44 66.88 64.33 49.57 48.5 59.5 55.83 46.13 41.07
P(true) 54.57 52.88 45.45 44.74 57.02 55.49 48.81 37.84 57.11 55.21 43.93 47.05 61.49 59.03 44.37 44.69
FActScore 56.76 53.85 40.25 40.01 54.51 53.2 38.45 36.49 62.11 58.64 53.52 47.27 58.82 57.47 49.48 42.74
TruthfulQA HalluGuard 72.1 68.76 60.09 52.01 69.59 68.36 58.52 52.65 77.05 73.79 63.62 62.26 74.26 72.76 57.39 64.07
Inside 70.42 68.76 60.09 52.01 62.1 59.78 51.07 51.38 62.53 60.99 52.3 49.35 70.89 64.44 56.61 56.01
MIND 59.45 56.79 45.22 43.71 60.56 58.55 47.49 49.63 59.2 57.98 47.23 41.79 62.81 61.5 52.56 46.37
Perplexity 50.57 47.87 40.64 35.63 55.07 52.26 44.43 42.79 60.8 59.69 47.33 41.62 55.29 52.46 43.95 43.92
LN-Entropy 58.04 56.99 41.94 47.21 56.12 54.01 47.06 38.4 59.67 56.25 41.99 41.25 60.76 58.21 46.24 42.64
Energy 55.02 53.31 38.78 45.16 54.42 51.85 36.21 42.57 58.93 55.25 50.76 41.72 64.15 61.32 51.78 50.02
Semantic Ent. 61.01 57.08 43.35 45.2 51.48 47.81 34.15 38.16 54.44 53.33 36.62 40.35 66.75 63.85 51.11 46.71
Lexical Sim. 52.54 50.56 39.94 33.42 59.74 55.72 49.89 46.81 66.16 64.05 54.08 51.65 55.24 51.36 46.39 39.57
SelfCheckGPT 56.04 54.48 43.78 44.38 58.93 56.47 47.65 39.02 61.14 58.91 42.97 47.01 55.86 54.95 41.08 37.35
RACE 53.02 50.33 41.7 33.81 62.95 67.89 54.61 51.93 71.06 68.49 60.4 57.44 55.75 52.62 46.5 43.19
P(true) 55.52 53.41 38.33 38.38 54.88 53.1 38.22 40.96 55.8 52.01 40.88 38.72 57.18 55.16 46.19 38.21
FActScore 53.82 51.42 41.33 35.2 54.57 51.26 42.51 35.52 53.97 50.2 42.97 36.16 62.31 60.23 45.06 49.9
Q2: How does HalluGuard perform across LLMs of different scales? We further investigate whether the effectiveness of HalluGuard depends on model scale, as smaller backbones are typically more prone to hallucination. Table 3 reports representative results on small(Llama2-7B, Llama3-8B), mid-sized(Llama2-13B), and large-scale(Llama2-70B) models using SQuAD, GSM8K, and HaluEval. Across all settings, HalluGuard consistently surpasses baselines, with the largest margins on smaller models—for instance,
Figure 2: Ablation results comparing individual terms with ground-truth trends on SQuAD (top) and Math-500 (bottom).
72.89% AUPRCr on HaluEval with Llama2-7B, more than 10% above the second best. Mid-sized models also exhibit clear gains (e.g., 79.01% AUROCr on GSM8K), while even large-scale models like Llama2-70B see steady improvements (e.g., 83.8% AUROCr on SQuAD). Overall, HalluGuard benefits most on small backbones while maintaining consistent advantages across scales.
Table 3: Performance comparison across backbone scales (small, mid-sized, and large) on three benchmarks: SQuAD, GSM8K, HaluEval. We highlight the first and second best results.
Llama2-7B Llama-3-8B Llama2-13B Llama2-70B
AUROCr
AUPRCr
[…]
AUROCr
AUPRCr
[…]
AUROCr
AUPRCr
[…]
AUROCr
AUPRCr
[…]
SQuAD HalluGuard 81.05 77.16 71.18 64.38 79.56 78.29 67.97 63.27 81.45 78.39 64.39 65.07 83.8 81.77 70.46 73.24
Inside 73.63 75.74 65.22 59.11 76.13 72.44 65.62 62.94 74.68 74.81 61.01 59.51 81.24 75.09 69.48 62.4
MIND 64.57 61.11 52.39 53.13 62.29 59.58 44.49 48.61 68.64 66.95 54.92 52.49 73.46 71.71 57.76 56.77
Perplexity 63.93 61.77 46.97 48.2 70.51 67.51 55.71 52,68 70.19 69.22 60.33 54.82 74.23 70.88 62.24 58.05
LN-Entropy 65.96 64.22 53.43 52.84 63.7 60.4 46.19 42.85 61.66 59.16 49.05 46.27 72.44 68.91 56.77 52.63
Energy 59.83 56.11 46.19 43.18 64.41 61.02 56.17 46.21 61.02 59.73 48.26 42.08 69.01 66.19 58.44 49.82
Semantic Ent. 60.29 57.73 43.63 48.83 66.52 62.62 52.37 52.7 70.58 67.22 53.31 52.94 72.01 68.51 56.49 50.9
Lexical Sim. 70.31 69.08 53.97 53.31 66.43 63.56 53.19 50.96 68.53 67.42 50.73 54.12 68.95 67.91 60.52 56.56
SelfCheckGPT 68.26 67.09 60.06 57.31 73.99 72.15 65.26 54.02 65.47 61.65 53.12 49.89 73.07 70.49 56.59 54.65
RACE 71.35 69.23 59.18 54.73 68.17 66.02 54.65 53.06 64.19 60.45 47.53 45.66 64.05 62.39 54.38 50.07
P(true) 62.55 61.09 46.84 52.32 67.42 63.94 55.35 47.52 71.56 68.4 57.51 45.66 66.81 62.71 57.43 46.85
FActScore 70.32 68.63 58.13 53.01 71.2 69.45 61.92 54.91 66.65 63.2 56.41 53.42 68.33 65.26 56.93 48.46
GSM8K HalluGuard 75.89 72.83 62.29 63.46 75.2 72.9 63.62 61.79 79.01 76.73 64.38 64.97 77.33 73.97 60.48 61.26
Inside 74.61 68.35 58.57 62.58 73.73 67.51 56.02 57.28 75.79 76.26 60.91 59.77 72.3 72.26 54.49 58.39
MIND 65.88 63.4 48.28 48.17 66.57 65.55 48.84 53.4 61.49 59.55 51.63 51.45 66.41 63.44 52.05 53.57
Perplexity 66.23 64.1 53.52 52.31 57.61 53.63 41.37 41.59 60.96 58.67 46.27 47.44 64.32 62.81 51.15 51.3
LN-Entropy 59.45 55.95 43.04 44.08 68.22 66.05 53.03 53.21 61.31 58.90 45.83 40.86 61.81 60.46 44.5 44.76
Energy 58.15 54.71 43.65 36.71 59.79 56.52 50.31 42.23 57.58 56.07 43.39 38.94 65.27 62.94 52.8 46.6
Semantic Ent. 57.95 54.68 42.78 41.95 66.9 64.81 50.47 55.36 62.72 59.09 49.33 44.35 60.63 57.01 46.22 40.24
Lexical Sim. 65.8 63.7 52.12 54.07 63.29 59.87 53.17 50.02 63.83 60.20 54.43 44.82 63.27 59.41 47.42 47.38
SelfCheckGPT 60.99 57.54 49.28 44.43 65.72 62.01 54.49 50.34 57.98 54.58 46.72 39.86 68.06 65.09 52.99 50.89
RACE 63.37 62.33 53.53 49.94 64.49 61.47 53.28 47.55 64.20 61.96 50.15 45.35 68.35 66.66 50.41 51.16
P(true) 65.95 63.63 54.95 48.25 62.59 58.88 47.21 42.2 67.08 65.60 53.66 55.12 60.16 58.14 47.73 49.49
FActScore 56.69 53.71 45.78 39.52 65.69 61.95 53.69 46.06 55.76 54.17 44.91 43.18 59.84 55.85 44.05 39.49
HaluEval HalluGuard 75.72 72.89 66.65 63.15 73.43 71.19 64.95 54.8 78.15 74.15 65.39 61.14 80.79 79.54 67.68 68.51
Inside 71.33 67.63 59.73 53.15 67.95 64.93 60.31 52.21 72.01 71.97 56.51 60.64 74.62 68.33 62.22 64.4
MIND 54.8 51.43 44.15 43.34 64.54 60.89 49.09 45.13 55.05 53.28 39.16 45.17 57.98 56.01 45.82 41.69
Perplexity 54.02 52.53 38.76 40.51 61.31 59.36 50.62 46.01 54.99 51.39 42.64 35.64 62.85 60.59 48.29 43.85
LN-Entropy 59.47 58.33 50.2 46.91 64.89 60.72 51.78 46.39 65.18 63.53 49.70 48.09 60.16 58.89 50.29 48.42
Energy 62.29 59.6 50.68 42.24 62.74 61.61 50.17 52.01 60.54 59.04 43.53 50.37 60.13 58.44 48.79 48.01
Semantic Ent. 59.39 55.94 48.53 46.35 55.25 53.05 44.5 44.35 59.44 57.72 45.38 40.77 61.57 57.99 49.07 45.39
Lexical Sim. 63.61 61.16 55.01 44.75 56.59 55.39 44.45 45.57 53.46 52.06 41.34 40.57 64.37 60.92 54.29 50.86
SelfCheckGPT 64.29 61.83 48.4 45.49 65.44 63.13 57.02 48.23 65.24 63.52 53.71 54.33 57.12 55.26 40.5 43.06
RACE 59.78 59.14 48.1 40.47 61.98 60.32 48.08 46.29 60.65 59.11 49.92 44.51 62.11 58.24 40.5 43.06
P(true) 57.46 54.8 41.84 40.47 56.32 54.04 42.55 43.75 65.77 63.01 49.98 45.47 55.75 54.94 44.14 43.97
FActScore 63.93 61.33 46.9 51.87 61.73 57.85 49.92 42.15 65.15 63.71 55.98 54.61 62.66 60.3 53.13 46.42
Q3: How does each term capture trends across task families? As shown in Figure 2, each term faithfully tracks the ground-truth trend within its respective task family. On data-centric SQuAD, the data-driven term closely follows the dashed gold curve across the variant hallucination rate, capturing the smooth AUROC decline. On reasoning-oriented MATH-500, the reasoning-driven term mirrors the monotonic AUROC drop as reasoning drift increases. These results show that each term is well matched to its task family and faithfully tracks performance trends as hallucination rates rise.
4.3 Test-Time Inference
Test-time reasoning remains challenging, as models need to generate coherent multi-step solutions without drifting into errors. To assess whether hallucination detection can mitigate this difficulty, we integrate detectors into beam search and evaluate Qwen2.5-Math-7B on MATH-500 and Llama3.1-8B on Natural. As shown in Table 4, HalluGuard achieves the strongest gains: on MATH-500, it reaches 81.00% accuracy, around 10% higher than IO Prompt; on Natural, it attains 70.96%, exceeding IO Prompt by 15.72%. These results demonstrate that HalluGuard not only detects hallucinations but also strengthens test-time reasoning by guiding models toward more reliable solutions.
Table 4: Performance of hallucination score-guided test-time inference across reasoning tasks. We highlight the first and second best results.
Dataset IO Prompt Ours Inside MIND Perplexity LN-Entropy Energy Semantic Ent. SelfCheck- GPT RACE P(true) FActScore
MATH-500 72.70 81.00 74.90 77.10 77.10 76.20 78.00 72.50 74.00 75.10 67.10 71.60
Natural 55.24 70.96 67.42 68.32 67.51 68.04 68.59 68.10 65.68 66.90 68.16 67.74
4.4 Case Study
Fine-grained hallucinations—lexically similar yet semantically incorrect outputs—pose a particular challenge for detection. To evaluate whether HalluGuard can comprehensively capture such subtle errors, we use the PAWS dataset(Zhang et al., 2019), which contrasts paraphrases with high surface overlap but divergent meanings. Following Li et al. (2025), we adopt ROUGE-based reference signals for evaluation (LABEL:tab:semantic_results). Across model scales, HalluGuard consistently surpasses baselines: it achieves 90.18% AUROC and 87.64% AUPRC on Llama2-70B, and 91.24% AUROC and 88.53% AUPRC on QwQ-32B—exceeding the next-best method by nearly five points. Even on GPT-2, it leads with 83.27% AUROC and 80.46% AUPRC. These results confirm HalluGuard’s effectiveness in capturing fine-grained semantic inconsistencies beyond benchmark settings.
Table 5: Results on PAWS measuring semantic hallucination detection with Llama-3.2-3B, Llama2-70B, and QwQ-32B. We highlight the first and second best results.
Method Ours Inside MIND Perplexity LN-Entropy Energy Semantic Ent. Lexical Sim. SelfCheck- GPT RACE P(true) FActScore
Llama3.2 AUROC 85.63 80.46 78.93 71.27 72.19 73.05 75.11 64.58 77.82 79.47 73.56 68.44
AUPRC 82.14 77.28 75.41 67.55 68.34 70.22 72.41 59.67 73.41 76.28 70.43 63.58
Llama2 AUROC 90.18 85.47 83.92 75.68 76.23 77.14 79.06 68.35 82.71 84.26 77.39 72.62
AUPRC 87.64 82.38 81.06 71.42 72.59 74.28 76.32 63.44 78.89 81.73 74.18 67.58
QwQ AUROC 91.24 85.41 84.56 76.72 77.43 78.29 80.42 69.54 83.59 86.38 78.53 73.46
AUPRC 88.53 82.27 81.37 72.63 73.29 75.44 77.18 64.27 79.42 83.41 75.21 68.32
5 Related Work
In this section, we review prior hallucination-detection methods by their detection target–Data-driven hallucinations and reasoning-driven hallucinations.
Detecting Data-Driven Hallucinations. Recent work has shown that internal activations encode rich indicators of such flaws. Chen et al. (2024a) proposed Eigenscore, which computes statistics of hidden representations from the eigen matrix to estimate hallucination risk. Su et al. (2024) introduced MIND, an unsupervised detector that models temporal dynamics of hidden states without requiring labels, along with HELM benchmark to enable standardized evaluation. Azaria and Mitchell (2023) demonstrated using linear probes on intermediate states to predict truthfulness.
Detecting Reasoning-Driven Hallucinations. There are other works targeting inference-time inconsistencies during generation—such as logical errors, instability across decoding steps, or temporal drift in extended outputs. Manakul et al. (2023) proposed SelfCheckGPT, which assesses self-consistency by sampling multiple candidate generations and measuring their alignment using entailment and lexical overlap. Kalai and Vempala (2024) introduced a suite of calibration-based uncertainty scores designed to capture hallucination risk directly from output distributions. Ding et al. (2025) proposed ReActScore, which integrates entropy with intermediate reasoning traces to detect failures in multi-step decision-making. FActScore(Min et al., 2023) decomposes outputs into atomic factual units and verifies each against retrieved passages using entailment-based scoring.
6 Conclusion
The reliability of LLMs is often undermined by hallucinations, which arise from two main sources: data-driven, caused by flawed knowledge acquired during training, and reasoning-driven, stemming from inference-time instabilities in multi-step generation. Although these hallucinations frequently evolve in practice, existing detectors usually target only one source and lack a solid theoretical foundation. To address this gap, we propose a unified theoretical framework–a Hallucination Risk Bound, which formally decomposes hallucination risk into data-driven and reasoning-driven components, offering a principled view of how hallucinations emerge and evolve during generation. Building on this foundation, we introduce HalluGuard, a NTK–based score that measures sensitivity to semantic perturbations and captures internal instabilities, thereby enabling holistic detection of both data-driven and reasoning-driven hallucinations. We evaluate HalluGuard across 10 diverse benchmarks, 11 competitive baselines, and 9 popular LLM backbones, where it consistently achieves state-of-the-art performance, demonstrating robustness and practical efficacy. Looking forward, leveraging HalluGuard’s sensitivity to error propagation offers a promising pathway for developing prognostic indicators in interactive multi-turn dialogues, enabling systems to predict and preempt hallucinations before they fully manifest.
Reproducibility Statement
We have taken several measures to ensure the reproducibility of our work. A complete description of the theoretical framework, including the formal assumptions and proofs of the Hallucination Risk Bound, is provided in Section 3 and Appendix A. Detailed experimental settings and evaluation protocols are documented in Section 4 and Section C.1, covering all 10 benchmarks, 11 baselines, and 9 LLM backbones. Together, these resources ensure that both our theoretical claims and empirical results can be independently validated and extended by the community.
Ethics Statement
This study is based exclusively on publicly available datasets and open-source large language models, and does not involve human subjects or the use of private data. All scientific concepts, methodological designs, experimental implementations, and resulting conclusions remain entirely the responsibility of the authors.
Appendix A Proof of Hallucination Risk Bound
A.1 Assumptions Validation
We provide theoretical and practical justification for the assumptions adopted in Section 3.2, which serve to ensure the well-posedness and interpretability of the proposed Hallucination Risk Bound. These assumptions follow standard practice in NTK-based analyses and stability theory, and are consistent with the empirical behavior observed in modern large language models.
Assumption A1 (Hilbert/RKHS structure with bounded second moment).
This assumption aligns with the classical Neural Tangent Kernel (NTK) approximation regime, where the model’s feature mapping is embedded in a reproducing kernel Hilbert space (RKHS) and the induced kernel admits a well-defined second moment. Such conditions are fundamental to the convergence and generalization analyses of infinitely wide neural networks, and are widely adopted in NTK theory (Jacot et al., 2020). In practice, bounded second-moment behavior is consistent with the hidden-state distributions observed across all evaluated LLMs, as reflected by stable activation statistics and NTK spectral profiles(Lee et al., 2020b).
Assumption A2 (Local Lipschitz continuity of the encoder $\Phi$).
This assumption reflects standard smoothness conditions in high-dimensional learning theory, ensuring that small perturbations in the input space induce controlled deviations in the encoded representation (Vershynin, 2018). Such local Lipschitz behavior is commonly invoked to guarantee stability under perturbations and is consistent with theoretical analyses of deep representations.
Assumption A3 (Local smoothness / second-order expansion).
This assumption corresponds to the classical NTK linearization framework, which approximates the behavior of wide neural networks through a local second-order expansion around a set of reference points (Lee et al., 2020a; Chizat et al., 2020). Importantly, our formulation requires this condition only locally around the $K$ sampled trajectories used by HalluGuard, rather than globally across the entire model parameter space. This localized validity preserves theoretical soundness while avoiding unrealistic global smoothness requirements that are known to be overly restrictive in large-scale models.
A.2 Bound Proof
We restate the main inequality from Section 3.2:
$\|u^{*}-u_{h}\|\leq\left[1+k_{\mathrm{pt}}\log\mathcal{O}(P,L)+k\frac{\epsilon_{\mathrm{mismatch}}}{\mathrm{Signal}_{k}}\right]\inf_{u\in U_{h}}\|u^{*}-u\|+\left|\mathcal{L}\right|\exp\left(-\frac{K\epsilon^{2}}{C}\right)\alpha\left(e^{\beta T}-1\right).$ (8)
Step 1: Triangle inequality split.
We define the hallucination decomposition by writing:
$\|u^{*}-u_{h}\|=\|u^{*}-\mathbb{E}[u_{h}]+\mathbb{E}[u_{h}]-u_{h}\|\leq\|u^{*}-\mathbb{E}[u_{h}]\|+\|u_{h}-\mathbb{E}[u_{h}]\|.$
We denote the first term as the deterministic approximation error (bias) and the second term as the stochastic residual (variance).
Step 2: Approximation term via Céa’s lemma.
Assume $\mathbb{E}[u_{h}]$ is the Galerkin projection of $u^{*}$ in a coercive bilinear form $a(\cdot,\cdot)$, i.e., for all $v\in U_{h}$,
$a(\mathbb{E}[u_{h}],v)=\ell(v).$
Then, by Céa’s lemma, we have:
$\|u^{*}-\mathbb{E}[u_{h}]\|\leq\frac{\Lambda}{\gamma}\inf_{u\in U_{h}}\|u^{*}-u\|,$
where $\Lambda$ and $\gamma$ are continuity and coercivity constants of $a(\cdot,\cdot)$, respectively.
Step 3: Variance term via Bernstein concentration.
Let $\ell_{h}:=\frac{1}{|\mathcal{L}|}\sum_{i=1}^{|\mathcal{L}|}\ell_{i}$ be the empirical supervision functional from finite labeled chains. Define the fluctuation:
$\Delta\ell:=\ell_{h}-\ell,$
and the residual:
$r:=u_{h}-\mathbb{E}[u_{h}],\quad\text{so that}\quad A_{h}r=\Delta\ell.$
Applying operator norm bounds and covering number uniformization (cf. Vershynin, 2018), we have with high probability:
$\|r\|\leq|\mathcal{L}|\exp\left(-\frac{K\epsilon^{2}}{C}\right)\alpha(e^{\beta T}-1),$
which completes the proof.
Step 4: Substitution.
Combining both terms yields:
$\|u^{*}-u_{h}\|\leq\frac{\Lambda}{\gamma}\inf_{u\in U_{h}}\|u^{*}-u\|+|\mathcal{L}|\exp\left(-\frac{K\epsilon^{2}}{C}\right)\alpha(e^{\beta T}-1).$
We now bound $\Lambda/\gamma$ via NTK decomposition.
A.3 Decomposition of NTK Continuity Constant
Let $a(\cdot,\cdot)$ denote the bilinear form induced by the NTK in the finite-width regime. We decompose:
$a=a_{0}+\delta_{\mathrm{pt}}+\delta_{\mathrm{mm}},$
where $a_{0}$ is the infinite-width baseline kernel, $\delta_{\mathrm{pt}}$ is the perturbation due to pre-training noise, and $\delta_{\mathrm{mm}}$ is the domain mismatch from fine-tuning. The continuity constant satisfies:
$\Lambda=\Lambda_{0}+\Delta_{\mathrm{pt}}+\Delta_{\mathrm{mm}}.$
Bounding $\Delta_{\mathrm{pt}}$.
Following Jacot et al. (2020), we apply matrix concentration to finite-width NTK:
$\Delta_{\mathrm{pt}}\leq\gamma k_{\mathrm{pt}}\log\mathcal{O}(P,L).$
Bounding $\Delta_{\mathrm{mm}}$.
Using spectral generalization bounds under data distribution shift (Lee et al., 2020b), we have:
$\Delta_{\mathrm{mm}}\leq\gamma k\frac{\epsilon_{\mathrm{mismatch}}}{\mathrm{Signal}_{k}}.$
Substituting both into the bound for $\Lambda/\gamma$, we get:
$\frac{\Lambda}{\gamma}\leq 1+k_{\mathrm{pt}}\log\mathcal{O}(P,L)+k\frac{\epsilon_{\mathrm{mismatch}}}{\mathrm{Signal}_{k}}.$
Appendix B HalluGuard Derivation and Interpretation
B.1 Preliminaries and Notation
Let $\mathcal{K}\in\mathbb{R}^{r\times r}$ be the NTK Gram matrix formed on $r$ light semantic perturbations (see Assumptions A1–A4 in the main theory section). Denote its eigen decomposition by $\mathcal{K}=V\Lambda V^{\top}$ with
$\Lambda=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{r}),\qquad\lambda_{1}\geq\cdots\geq\lambda_{r}>0.$
Let $\lambda_{\max}:=\lambda_{1}$, $\lambda_{\min}:=\lambda_{r}$, $\kappa(\mathcal{K}):=\lambda_{\max}/\lambda_{\min}$, and $\det(\mathcal{K})=\prod_{i=1}^{r}\lambda_{i}$. Let $\Phi$ denote the NTK feature matrix whose columns span the hypothesis subspace $U_{h}$, so that $\mathcal{K}=\Phi^{\top}\Phi$, $\|\Phi\|_{2}=\sqrt{\lambda_{\max}}$, and $\sigma_{\min}(\Phi)=\sqrt{\lambda_{\min}}$. For the autoregressive decoder, let $J_{t}$ be the step-$t$ input–output Jacobian, and write $\sigma_{\max}:=\sup_{t}\|J_{t}\|_{2}$.
We will use the following two standard inequalities repeatedly:
$Maclaurin/AM--GMoneigenvalues:$ $\Big(\prod_{i=1}^{r}\lambda_{i}\Big)^{\!\!1/r}\;\leq\;\frac{1}{r}\sum_{i=1}^{r}\lambda_{i}\;=\;\frac{\mathrm{tr}(\mathcal{K})}{r},$ (9)
$Submultiplicativity:$ $\|AB\|_{2}\;\leq\;\|A\|_{2}\,\|B\|_{2}.$ (10)
B.2 Representational Adequacy via $\det(\mathcal{K})$ with Explicit Constants
Assumptions for this subsection.
Beyond A1–A3, we assume a mild source condition and a spectral envelope:
  • S1
    (Source condition) There exist $s>0$ and $R_{s}>0$ such that $u^{*}\in\mathrm{Range}(\Lambda^{s})$, i.e., $\sum_{i=1}^{r}\frac{\langle u^{*},v_{i}\rangle^{2}}{\lambda_{i}^{2s}}\leq R_{s}^{2}$. This is standard in kernel approximation and encodes RKHS regularity.
  • S2
    (Spectral envelope) There exist constants $0<\underline{\lambda}\leq\overline{\lambda}<\infty$ and $\alpha>1$ such that $\lambda_{i}\leq\overline{\lambda}$ for all $i$ and $\lambda_{r}\geq\underline{\lambda}\,r^{-\alpha}$. (Polynomial decay is a common stylization; other envelopes can be treated similarly.)
Lemma B.1 (Best-approximation error under source condition).
Let $U_{h}=\mathrm{span}\{v_{1},\ldots,v_{r}\}$. Under S1,
$\inf_{u\in U_{h}}\|u^{*}-u\|\;=\;\|u^{*}-\Pi_{U_{h}}u^{*}\|\;\leq\;R_{s}\,\lambda_{r+1}^{\,s},$
where $\lambda_{r+1}$ denotes the next-eigenvalue of the infinite-dimensional kernel operator (or, equivalently, the empirical tail eigenvalue if more perturbations are added).
Proof.
Write $u^{*}=\sum_{i\geq 1}c_{i}v_{i}$ with $c_{i}=\langle u^{*},v_{i}\rangle$. Then $\|u^{*}-\Pi_{U_{h}}u^{*}\|^{2}=\sum_{i>r}c_{i}^{2}\leq\sum_{i>r}\lambda_{i}^{2s}\cdot\frac{c_{i}^{2}}{\lambda_{i}^{2s}}\leq\lambda_{r+1}^{2s}\sum_{i>r}\frac{c_{i}^{2}}{\lambda_{i}^{2s}}\leq\lambda_{r+1}^{2s}R_{s}^{2}.$
To connect $\lambda_{r+1}$ (or $\lambda_{r}$) to $\det(\mathcal{K})$, we need an explicit lower bound of the form $\lambda_{r}\geq\underline{c}\,\det(\mathcal{K})^{\,\theta}$ with constants $(\underline{c},\theta)$ depending on the spectral envelope. The following inequality suffices.
Lemma B.2 (Lower-bounding $\lambda_{r}$ by $\det(\mathcal{K})$).
Suppose $\lambda_{i}\leq\overline{\lambda}$ for all $i$ and $\lambda_{r}>0$. Then
$\lambda_{r}\;\geq\;\frac{\det(\mathcal{K})}{\overline{\lambda}^{\,r-1}}\qquad\text{and}\qquad\lambda_{r}^{\,s}\;\geq\;\frac{\det(\mathcal{K})^{\,s}}{\overline{\lambda}^{\,s(r-1)}}.$
Proof.
Since $\det(\mathcal{K})=\prod_{i=1}^{r}\lambda_{i}\leq\overline{\lambda}^{\,r-1}\lambda_{r}$, we obtain $\lambda_{r}\geq\det(\mathcal{K})/\overline{\lambda}^{\,r-1}$. Raising to power $s$ yields the second inequality. ∎
Theorem B.3 (Determinant-based adequacy bound with explicit constants).
Under A1–A3 and S1S2,
$\inf_{u\in U_{h}}\|u^{*}-u\|\;\leq\;C_{d}\;\det(\mathcal{K})^{-\,c_{d}}\;\|u^{*}\|,$
with
$c_{d}\;=\;\frac{s}{\,r-1\,}\quad\text{and}\quad C_{d}\;=\;\overline{\lambda}^{\,s}\,\frac{R_{s}}{\|u^{*}\|}.$
Moreover, if the empirical spectrum satisfies $\lambda_{r}\geq\underline{\lambda}\,r^{-\alpha}$, one may choose
$c_{d}\;=\;\min\!\left\{\frac{s}{\,r-1\,},\;\frac{s}{\alpha}\cdot\frac{1}{\log\!\big(\tfrac{\overline{\lambda}^{r}}{\det(\mathcal{K})}\big)}\right\},$
which improves with slower decay (smaller $\alpha$).
Proof.
By Lemma B.1 with $\lambda_{r+1}\leq\lambda_{r}$, $\inf_{u\in U_{h}}\|u^{*}-u\|\leq R_{s}\,\lambda_{r}^{\,s}.$ Lemma B.2 gives $\lambda_{r}^{\,s}\geq\det(\mathcal{K})^{\,s}/\overline{\lambda}^{\,s(r-1)}$; rearranging,
$\inf_{u\in U_{h}}\|u^{*}-u\|\;\leq\;R_{s}\,\overline{\lambda}^{\,s(r-1)}\,\det(\mathcal{K})^{-\,s}.$
Rescale constants relative to $\|u^{*}\|$ by setting $C_{d}:=\overline{\lambda}^{\,s}\,(R_{s}/\|u^{*}\|)$ and $c_{d}:=s/(r-1)$ to obtain the stated form:
$\inf_{u\in U_{h}}\|u^{*}-u\|\;\leq\;\big(\overline{\lambda}^{\,s}\,\tfrac{R_{s}}{\|u^{*}\|}\big)\;\det(\mathcal{K})^{-\,s/(r-1)}\;\|u^{*}\|.$
The variant using the envelope $\lambda_{r}\geq\underline{\lambda}\,r^{-\alpha}$ is obtained by combining $\det(\mathcal{K})\leq\overline{\lambda}^{\,r-1}\lambda_{r}$ with the explicit lower bound on $\lambda_{r}$, yielding the alternative exponent shown. ∎
Numerical note (stable surrogate).
In practice we use $\log\det(\mathcal{K})$ via Cholesky and aggregate with $z$-normalization across components to avoid scale domination by any single term.
B.3 Rollout Amplification via Jacobian Products (Exact Constants)
Theorem B.4 (Amplification bound with exact constant).
Let $J_{t}$ be the step-$t$ Jacobian and $\sigma_{\max}:=\sup_{t}\|J_{t}\|_{2}$. Then
$\Big\|\prod_{t=1}^{T}J_{t}\Big\|_{2}\;\leq\;\prod_{t=1}^{T}\|J_{t}\|_{2}\;\leq\;\sigma_{\max}^{\,T}.$
Defining $\beta:=\log\sigma_{\max}$ gives $e^{\beta T}=\sigma_{\max}^{T}$, hence
$e^{\beta T}\;\leq\;\sigma_{\max}^{T},$
with equality if and only if $\|J_{t}\|_{2}=\sigma_{\max}$ for all $t$ and the top singular directions align across factors.
Proof.
The first inequality is equation 10 applied iteratively. The second is by definition of $\sigma_{\max}$. Setting $\beta=\log\sigma_{\max}$ yields equality in the worst case. Alignment of top singular vectors is the tightness condition for submultiplicativity. ∎
Token-dependent refinement.
If one defines $\sigma_{t}:=\|J_{t}\|_{2}$ and $\beta_{\mathrm{avg}}:=\tfrac{1}{T}\sum_{t=1}^{T}\log\sigma_{t}$, then $\big\|\prod_{t=1}^{T}J_{t}\big\|_{2}\leq\exp\!\big(\sum_{t}\log\sigma_{t}\big)=e^{\beta_{\mathrm{avg}}T}$, which is tighter but requires per-step measurements.
B.4 Conditioning-Induced Variance with $\kappa(\mathcal{K})^{2}$ Scaling
We now give an explicit projector-perturbation derivation showing the quadratic dependence on the condition number.
Setup.
Let $P:=\Phi(\Phi^{\top}\Phi)^{\dagger}\Phi^{\top}$ be the orthogonal projector onto $U_{h}$; then the linearized output is $u_{h}=Pu^{*}$. Consider a feature perturbation $\Delta\Phi$ induced by a prefix perturbation $\delta$ satisfying
$\|\Delta\Phi\|_{2}\;\leq\;L_{\Phi}\,\|\delta\|\quad\text{(A2/A3)}.$
Let the perturbed projector be $\widetilde{P}:=(\Phi+\Delta\Phi)\big((\Phi+\Delta\Phi)^{\top}(\Phi+\Delta\Phi)\big)^{\dagger}(\Phi+\Delta\Phi)^{\top}$ and define $\Delta P:=\widetilde{P}-P$.
Lemma B.5 (Projector perturbation bound).
There exists an absolute constant $C_{\Pi}>0$ such that
$\|\Delta P\|_{2}\;\leq\;C_{\Pi}\;\frac{\|\Phi\|_{2}}{\sigma_{\min}(\Phi)^{2}}\;\|\Delta\Phi\|_{2}\;=\;C_{\Pi}\;\frac{\sqrt{\lambda_{\max}}}{\lambda_{\min}}\;\|\Delta\Phi\|_{2}\;=\;C_{\Pi}\;\kappa(\mathcal{K})\,\frac{\|\Delta\Phi\|_{2}}{\sqrt{\lambda_{\min}}}.$
Proof idea.
[…]
$\|\Delta P\|_{2}\;\leq\;2\,\|(\Phi^{\top}\Phi)^{\dagger}\|_{2}\,\|\Phi^{\top}\Delta\Phi\|_{2}\;+\;\mathcal{O}(\|\Delta\Phi\|_{2}^{2}).$
Since $\|(\Phi^{\top}\Phi)^{\dagger}\|_{2}=1/\lambda_{\min}$ and $\|\Phi^{\top}\Delta\Phi\|_{2}\leq\|\Phi\|_{2}\,\|\Delta\Phi\|_{2}=\sqrt{\lambda_{\max}}\|\Delta\Phi\|_{2}$, the result follows for sufficiently small $\|\Delta\Phi\|_{2}$, absorbing lower-order terms into $C_{\Pi}$. ∎
Theorem B.6 (Variance amplification with explicit constant).
Let $u_{h}(\Phi)=Pu^{*}$ and $u_{h}(\Phi+\Delta\Phi)=\widetilde{P}u^{*}$. Then
$\|u_{h}(\Phi+\Delta\Phi)-u_{h}(\Phi)\|\;\leq\;C_{\Pi}\,\kappa(\mathcal{K})\,\frac{\|\Delta\Phi\|_{2}}{\sqrt{\lambda_{\min}}}\,\|u^{*}\|.$
If $\Delta\Phi$ is induced by a random prefix perturbation $\delta$ with $\|\Delta\Phi\|_{2}\leq L_{\Phi}\|\delta\|$ and $\mathbb{E}\|\delta\|^{2}=\sigma_{\delta}^{2}$, then
$\mathrm{Var}[u_{h}]\;\leq\;\mathbb{E}\|u_{h}(\Phi+\Delta\Phi)-u_{h}(\Phi)\|^{2}\;\leq\;c_{v}\;\kappa(\mathcal{K})^{2}\;\|\delta\|^{2},$
with
$c_{v}\;=\;C_{\Pi}^{2}\,\frac{L_{\Phi}^{2}\,\|u^{*}\|^{2}}{\lambda_{\min}}.$
Proof.
By Lemma B.5, $\|u_{h}(\Phi+\Delta\Phi)-u_{h}(\Phi)\|=\|\Delta P\,u^{*}\|\leq\|\Delta P\|_{2}\|u^{*}\|\leq C_{\Pi}\,\kappa(\mathcal{K})\,\frac{\|\Delta\Phi\|_{2}}{\sqrt{\lambda_{\min}}}\,\|u^{*}\|.$ Square both sides and take expectation over $\delta$, using $\|\Delta\Phi\|_{2}\leq L_{\Phi}\|\delta\|$, to obtain the stated variance bound with the explicit constant $c_{v}$. ∎
Interpretation.
The $\kappa(\mathcal{K})^{2}$ factor arises from two sources: (i) $\kappa(\mathcal{K})$ from the projector sensitivity (Lemma B.5), and (ii) $1/\lambda_{\min}$ from converting $\|\Delta P\|_{2}$ to a mean-squared bound after squaring and averaging, yielding an overall $\kappa^{2}$-scaling in the variance constant.
B.5 Consolidation: Compact Surrogate Consistent with the Risk Decomposition
Combining Theorem B.3, Theorem B.4, and Theorem B.6, we obtain a computable surrogate aligned with the Hallucination Risk Bound:
$\text{Adequacy: }\det(\mathcal{K})\quad\;\;\text{Amplification: }\log\sigma_{\max}\quad\;\;\text{Conditioning penalty: }-\log\kappa(\mathcal{K})^{2}.$
This motivates the score
$\boxed{\;\text{{{HalluGuard}}}(u_{h})\;=\;\det(\mathcal{K})\;+\;\log\sigma_{\max}\;-\;\log\kappa(\mathcal{K})^{2}\;}$
with the following explicit, implementation-ready notes:
  • Use $\log\det(\mathcal{K})$ via Cholesky for stability; replace $\det$ in the score with $\log\det$ if desired (monotone equivalent).
  • Estimate $\sigma_{\max}$ either as $\sup_{t}\|J_{t}\|_{2}$ or its tighter average form $\beta_{\mathrm{avg}}=\tfrac{1}{T}\sum_{t}\log\|J_{t}\|_{2}$ (then use $\beta_{\mathrm{avg}}$ in place of $\log\sigma_{\max}$).
  • $z$-normalize each component across a validation set before summation to avoid scale dominance; optionally fit task-specific weights if permitted.
Appendix C Experiment
C.1 Setup
Implementation Framework.
All experiments use PyTorch and HuggingFace Transformers with a fixed random seed for reproducibility. Unless otherwise noted, computations run in mixed precision (fp16). Hardware details (A100/H200) are reported once in the main setup section.
Generation Configuration.
For default evaluation of detectors, we use nucleus sampling with temperature =0.5=0.5, top-p =0.95=0.95, and top-k =10=10, decoding K=10K{=}10 candidate responses per input (unless otherwise specified). These decoding trajectories also operationalize semantic perturbations as natural variations within the model’s local predictive distribution, thereby instantiating a semantically proximate neighborhood around the primary response and capturing the local geometry of the reasoning manifold required for NTK construction. For score-guided test-time inference (Section 4.3), we use beam search (beam size =10=10) and score candidate trajectories at each step with the chosen detector. For stability analysis, HalluGuard extracts sentence representations from the final token at the middle transformer layer ($L/2$), which empirically preserves semantics relevant to truthfulness.
NTK-Based Score Computation.
For each set of generations, we form a task-specific NTK feature matrix and compute the semantic stability score from its eigenspectrum. We add a small ridge $\alpha=10^{-3}$ for numerical stability and compute singular values via SVD.
Perturbation Regularization.
To prevent pathological activations that amplify instability, HalluGuard clips hidden features using an adaptive scheme. We maintain a memory bank of N=3000N{=}3000 token embeddings and set thresholds at the top and bottom $0.2\%$ percentiles of neuron activations; out-of-range values are truncated to attenuate overconfident hallucinations.
Optimization.
Backbone language models are not fine-tuned. We train only HalluGuard’s lightweight projection layers using AdamW with learning rate selected from $\{1\times 10^{-5},\,5\times 10^{-5},\,1\times 10^{-4}\}$ and weight decay from $\{0.0,\,0.01\}$. The best setting is chosen on a held-out validation split.
Implementation Details.
For score-guided inference we apply beam search with beam size $10$, rescoring candidates stepwise with different hallucination detectors.
Ablation Setup.
[…]
C.2 Ablation Study on $\log\kappa^{2}$
To empirically validate the necessity of the stability term $\log\kappa^{2}$, we performed a controlled ablation on MATH-500. We systematized the reasoning drift ($d$) by progressively increasing the perturbation budget $\rho$ and rollout horizon $T$. As shown in Figure 3, the absence of this term leads to severe instability. While the ablated model (orange dashed line) performs competitively in low-drift regimes ($d<0.15$), it exhibits significant performance volatility as the reasoning task becomes more complex. In contrast, the full HalluGuard score (green solid line) effectively penalizes these ill-conditioned regimes, maintaining a smooth and robust detection profile. This confirms that $\log\kappa^{2}$ functions as an essential spectral regularizer, preventing the score from becoming unreliable under high-entropy inference states.
Figure 3: Ablation study of the stability term ($\log\kappa^{2}$) on MATH500.
C.3 Computational Efficiency Analysis
To assess practical deployment feasibility, we measured inference latency on an NVIDIA A100/H200 GPU. Our setup utilizes batched parallel sampling to generate $K=10$ trajectories, ensuring sub-linear scaling of the computational cost. The core HalluGuard operations—specifically feature clipping and computing the NTK score via the Gram matrix—add minimal latency, requiring less than 1 ms of post-processing time per query.
Figure 4: Per-Question Inference Time (Seconds) on BBH Across Hallucination Detection Methods.
Figure 5: Per-Question Inference Time (Seconds) on HaluEval Across Hallucination Detection Methods.
Figure 6: Per-Question Inference Time (Seconds) on Math500 Across Hallucination Detection Methods.
Figure 7: Per-Question Inference Time (Seconds) on RAGTruth Across Hallucination Detection Methods.
Figure 8: Per-Question Inference Time (Seconds) on SQuaD Across Hallucination Detection Methods.
Figure 9: Per-Question Inference Time (Seconds) on TruthfulQA Across Hallucination Detection Methods.
C.4 Detection Performance Analysis
Across all five model families and three benchmark regimes, HalluGuard consistently achieves state-of-the-art detection performance, particularly in the safety-critical low-FPR regions as shown in Table 6.
Table 6: Performance comparison on representative benchmarks: data-centric (RAGTruth), reasoning-oriented (BBH), and instruction-following (TruthfulQA).
GPT2 OPT-6.7B Mistral-7B QwQ-32B LLaMA2-13B
F1
TPR@10%
TPR@5%
F1
TPR@10%
TPR@5%
F1
TPR@10%
TPR@5%
F1
TPR@10%
TPR@5%
F1
TPR@10%
TPR@5%
RAGTruth HalluGuard 81.22 74.86 61.41 77.03 73.52 59.12 83.19 79.44 69.21 85.91 80.13 63.52 74.66 68.91 57.42
Inside 66.12 59.72 48.31 72.91 70.25 60.37 70.45 68.12 52.41 79.03 74.66 61.09 73.08 70.11 55.26
MIND 58.33 54.11 38.72 62.55 57.81 47.65 71.91 66.74 54.39 64.02 59.12 45.63 68.55 63.50 48.78
Perplexity 55.42 51.20 40.51 63.72 60.13 49.14 69.74 66.51 52.18 70.42 65.41 55.32 60.18 57.01 44.75
LN-Entropy 62.17 57.52 46.44 58.33 52.99 43.28 65.30 61.27 49.92 67.15 62.42 51.33 63.28 59.07 46.14
Energy 59.71 56.23 44.81 60.44 57.18 45.03 63.54 59.42 48.62 72.09 68.15 58.42 66.10 61.33 49.41
Semantic Ent. 57.28 53.42 41.92 69.61 64.81 52.01 67.10 62.44 50.66 66.12 62.15 49.31 64.55 60.18 47.75
Lexical Sim. 61.41 57.09 45.03 65.81 61.44 49.51 62.50 59.12 50.92 70.91 67.53 55.21 66.29 59.88 51.03
SelfCheckGPT 56.22 52.84 40.63 60.79 55.68 45.72 63.12 59.47 48.33 66.54 62.92 51.41 68.21 65.12 53.60
RACE 60.12 56.50 44.90 64.12 59.77 49.22 65.44 61.55 52.73 69.61 66.31 53.92 62.55 59.42 45.66
P(true) 58.91 55.47 42.13 67.44 63.20 51.43 71.22 66.91 54.10 63.44 60.33 49.27 70.18 65.77 52.78
FActScore 62.10 58.21 46.33 59.22 54.14 44.32 63.87 60.77 47.98 68.33 64.02 53.41 65.92 61.37 49.84
BBH HalluGuard 78.33 74.11 65.42 74.91 69.14 62.10 80.22 76.88 68.21 82.55 78.91 70.45 79.10 74.25 67.92
Inside 65.41 61.22 52.83 71.02 67.10 60.21 68.17 64.75 53.92 79.17 72.33 64.22 67.10 63.52 55.91
MIND 54.12 50.22 40.11 57.21 53.44 41.52 63.92 59.88 47.01 61.55 57.14 48.83 65.11 60.22 49.52
Perplexity 52.91 49.33 40.44 61.88 58.12 49.22 62.91 59.42 50.11 59.91 55.72 49.03 60.88 57.41 48.62
LN-Entropy 59.12 55.44 44.92 54.61 51.75 43.18 66.44 63.21 54.09 62.75 59.12 47.52 68.20 64.88 55.41
Energy 53.94 51.22 45.03 56.12 52.14 44.61 64.55 60.11 49.99 68.21 65.12 52.84 66.41 62.77 50.22
Semantic Ent. 57.41 54.32 47.21 61.22 58.42 49.74 63.21 59.10 48.62 63.55 60.24 48.88 64.91 61.44 50.72
Lexical Sim. 50.41 46.77 38.92 60.71 57.11 45.55 59.42 56.88 48.91 70.33 67.10 55.32 58.33 55.42 47.41
SelfCheckGPT 55.21 52.14 43.92 58.10 55.78 46.22 62.82 59.90 50.44 65.22 62.44 54.21 63.44 60.77 52.33
RACE 56.14 53.72 43.88 63.11 59.71 52.81 65.77 62.55 50.72 58.88 55.14 46.18 66.10 62.41 49.81
P(true) 54.31 52.22 44.10 58.22 56.10 48.52 56.91 53.55 43.92 61.40 58.21 46.77 57.33 54.88 45.91
FActScore 56.20 52.42 41.77 55.44 52.12 41.14 61.62 58.22 51.33 59.33 56.42 49.14 63.44 60.22 52.44
TruthfulQA HalluGuard 75.11 71.20 63.21 70.44 67.55 58.12 78.92 74.22 65.33 76.44 72.01 59.92 79.33 75.11 66.08
Inside 71.10 68.55 60.77 61.77 59.44 50.10 63.88 61.33 53.41 69.22 65.10 55.14 62.14 59.94 52.80
MIND 57.44 54.91 45.33 59.92 56.88 48.33 58.72 56.14 47.21 61.21 58.88 52.02 60.44 58.20 49.03
Perplexity 49.52 46.71 38.84 54.12 51.74 43.90 59.72 57.55 46.88 54.44 51.72 42.55 60.33 57.21 47.41
LN-Entropy 57.11 54.88 42.98 55.33 52.41 45.91 59.66 56.22 43.10 60.44 58.02 46.22 61.41 57.17 43.88
Energy 54.11 52.17 38.91 53.44 51.14 36.88 58.21 54.77 49.92 63.02 60.44 51.33 58.41 55.33 50.42
Semantic Ent. 60.08 56.44 44.15 50.14 47.33 35.92 53.74 52.11 37.02 65.33 63.20 50.77 55.02 53.11 38.44
Lexical Sim. 51.22 49.20 39.03 58.72 54.71 48.77 65.71 63.50 53.10 54.77 51.44 45.88 66.41 64.14 54.88
SelfCheckGPT 55.72 53.44 42.78 58.33 55.72 47.14 60.88 57.44 43.91 55.42 54.44 40.77 61.72 59.51 44.10
RACE 52.22 49.88 41.44 63.14 66.88 54.05 70.55 67.11 59.77 55.44 52.11 45.33 71.33 68.22 60.02
P(true) 55.54 52.11 38.82 55.72 52.33 39.22 57.41 53.10 41.22 56.88 54.77 45.55 57.12 53.33 41.88
FActScore 52.91 50.14 40.44 54.11 50.22 41.33 52.88 49.91 42.55 61.55 59.22 44.72 53.41 50.71 43.10
We additionally expanded our evaluation to include SAPLMA, LLM-Check, and ITI. As shown in  Table 7, HalluGuard delivers the strongest performance not only on AUROC/AUPRC but also on deployment-critical, low-FPR operating points, including F1 and TPR at 5% and 10% FPR. Across all three benchmarks (RAGTruth, GSM8K, HaluEval) and all backbones (GPT-2 through QwQ-32B and LLaMA2-13B), HalluGuard consistently achieves the highest F1 and the highest or near-highest TPR under fixed low-FPR constraints. In contrast, SAPLMA and LLM-Check exhibit noticeably lower recall in the stringent 5% FPR regime. These results demonstrate that HalluGuard is better aligned with maintaining high detection sensitivity under tight false-positive budgets, a requirement that is central to reliable hallucination detection in real-world systems.
Table 7: Comparison with SAPLMA, LLM-Check and ITI across benchmarks and backbones.
Benchmark Method GPT2 OPT-6.7B Mistral-7B QwQ-32B LLaMA2-13B
AUROC AUPRC F1 TPR@10% TPR@5% AUROC AUPRC F1 TPR@10% TPR@5% AUROC AUPRC F1 TPR@10% TPR@5% AUROC AUPRC F1 TPR@10% TPR@5% AUROC AUPRC F1 TPR@10% TPR@5%
RAGTruth HalluGuard 75.51 73.40 81.22 74.86 61.41 80.13 76.77 77.03 73.52 59.12 82.31 80.79 83.19 79.44 69.21 84.59 81.15 85.91 80.13 63.52 77.51 75.30 74.66 68.91 57.42
RAGTruth SAPLMA 72.80 70.10 72.20 63.50 55.10 78.90 74.20 74.10 68.00 58.20 79.40 77.30 79.00 72.10 60.50 81.00 78.20 79.44 72.80 61.30 74.20 72.10 70.50 61.80 55.90
RAGTruth LLM-Check 68.10 64.50 63.90 55.20 44.80 72.30 68.40 66.50 57.90 46.30 75.20 71.60 67.40 60.30 48.70 76.10 73.20 68.90 61.10 49.50 71.60 68.90 63.20 55.40 46.10
RAGTruth ITI 69.30 65.80 66.10 57.90 47.90 73.10 69.20 68.20 59.80 49.10 76.00 72.50 69.40 61.80 50.90 77.20 74.10 70.50 62.40 51.70 72.80 70.10 65.40 57.10 47.80
GSM8K HalluGuard 72.04 69.88 78.33 74.11 65.42 72.57 70.31 74.91 69.14 62.10 80.62 77.30 80.22 76.88 68.21 75.81 74.68 82.55 78.91 70.45 79.01 76.73 79.10 74.25 67.92
GSM8K SAPLMA 69.20 66.10 70.10 62.00 54.40 70.80 67.20 71.80 64.10 56.30 77.10 74.00 76.20 69.50 59.80 73.90 71.20 76.50 70.10 60.70 75.40 72.30 74.00 67.10 59.10
GSM8K LLM-Check 65.40 61.50 62.40 54.10 46.20 68.10 64.30 67.50 59.20 49.80 73.40 69.80 64.90 57.90 48.30 71.20 67.90 67.80 60.30 50.40 72.10 68.50 64.20 56.60 48.00
GSM8K ITI 66.80 63.00 64.50 56.20 48.70 69.00 65.40 69.20 61.50 51.90 74.20 70.60 67.10 60.80 50.10 72.50 69.20 69.40 62.50 52.30 73.00 69.10 66.10 58.40 49.50
HaluEval HalluGuard 70.42 67.71 75.11 71.20 63.21 71.62 67.88 70.44 67.55 58.12 74.91 72.74 78.92 74.22 65.33 73.93 70.87 76.44 72.01 59.92 78.15 74.15 79.33 75.11 66.08
HaluEval SAPLMA 67.10 63.20 69.20 62.10 54.00 69.50 65.70 68.30 61.60 53.20 72.00 68.40 75.10 69.30 58.90 71.20 68.10 75.40 70.30 58.50 76.10 72.20 76.80 70.60 60.90
HaluEval LLM-Check 63.50 59.40 61.10 53.00 44.50 66.80 62.90 65.40 57.50 47.50 70.10 66.30 63.80 57.20 47.10 69.30 65.40 66.20 59.50 49.00 71.50 67.60 63.50 55.90 47.40
HaluEval ITI 64.80 60.70 63.40 55.20 46.80 67.40 63.50 66.90 58.60 49.40 71.00 67.20 66.10 59.10 48.60 70.20 66.30 68.10 61.10 50.60 72.30 68.20 65.20 57.50 48.70
C.5 Tightness of Bound
Evaluation of bound tightness.
To rigorously stress-test the Hallucination Risk Bound of Theorem 3.2, we conducted a controlled synthetic study grounded in the empirical reasoning-depth distribution of the Snowballing dataset (Zhang et al., 2023). We instantiated empirical hallucination trajectories by injecting low-variance Gaussian noise into the base components $\mathsf{D}(T)$ and $\mathsf{R}(T)$, comparing them against the closed-form theoretical prediction. As illustrated in Figure 10, while the theoretical curve acts as a conservative upper envelope, it exhibits a nearly parallel growth trajectory to the empirical risk. Crucially, it faithfully captures the exponential curvature and compounding dynamics of the Snowballing Effect. This confirms that the bound possesses high structural fidelity: it correctly models the scaling law of error propagation across depth ranges, validating its effectiveness as a ranking proxy despite the absolute numerical offset.
Figure 10: Empirical hallucination risk versus our theoretical bound
Evaluation of NTK proxy tightness.
To quantitatively validate that our NTK-based proxy faithfully captures the amplification behavior of stepwise Jacobians, we conduct a diagnostic experiment on GPT-2-small (117M), where per-step Jacobian norms are fully tractable. For a held-out set of GSM8K prompts and decoding steps $t\leq 18$, we compute:
  • the empirical stepwise Jacobian magnitude $\|J_{t}\|_{2}$, obtained via automatic differentiation on the next-token logits, and
  • our reasoning-driven NTK proxy, $\log\sigma_{\max}-\log\kappa^{2}$, as defined in Eq. (7), which upper-bounds the per-step amplification rate and penalizes spectral ill-conditioning of the NTK Gram matrix.
Figure 11 reports the scatter plot comparing the NTK proxy against empirical $\|J_{t}\|_{2}$ across all prompts and steps.
Figure 11: The NTK proxy closely tracks empirical Jacobian amplification on GPT-2-small, showing near-perfect monotonic alignment and a consistent conservative envelope across decoding depth.
Validation of Term Decomposition
To validate the architectural premise of our Hallucination Risk Bound Section 3.2, we visualize the evolution of the decomposed risk components across reasoning depth $T$ on the Snowballing dataset (Zhang et al., 2023). As shown in Figure Figure 12, the total risk is driven by two distinct dynamic behaviors. The data-driven term (green dotted line) exhibits linear or near-constant progression, reflecting static retrieval or knowledge-encoding errors that persist regardless of depth. In contrast, the reasoning-driven term (purple dotted line) demonstrates exponential amplification consistent with the Snowballing Effect, remaining negligible at shallow depths but rapidly dominating the total risk as $T$ increases.Crucially, this reveals a phase transition in hallucination dynamics: at lower depths ($T<15$), errors are primarily data-driven, whereas at higher depths, reasoning instability becomes the governing factor. This dichotomy empirically justifies our hybrid scoring mechanism, confirming that a unified detector must account for both the static semantic bias and the dynamic rollout instability to be effective across varying generation lengths.
Figure 12: Risk decomposition across reasoning depth T on Snowballing dataset.
C.6 Correlation of reasoning-driven and data-driven terms with different types of datasets
To empirically verify the independence of the proposed risk components, we analyzed their correlation with detection performance across distinct task families. As illustrated in Figure 14 and Figure 13, we observe a sharp geometric decoupling: the data-driven term aligns strongly with data-centric benchmarks (e.g., RAGTruth) while showing negligible correlation with reasoning tasks. Conversely, the reasoning-driven term dominates on reasoning-oriented datasets (e.g., MATH-500). This double dissociation reinforces the structural validity and orthogonality of our decomposition, confirming that each term captures a distinct, non-redundant failure mode.
Figure 13: Correlation Between data-driven and reasoning-driven terms and AUROC on Reasoning-Centric MATH500.
Figure 14: Correlation Between data-driven and reasoning-driven terms and AUROC on Data-Centric RAGTruth.
C.7 Case Study
Case Study 1 — GSM8K (Multi-step Arithmetic): Bias $\rightarrow$ Drift $\rightarrow$ Snowballing.
Task: “John saves \$3/day for four weeks and buys a \$12 toy. How much money does he have left?”
Ground truth: \$72.
Length (T) Model Behavior HalluGuard Response
T=1–8 Stable setup Correct restatement and arithmetic planning Data-driven term dominant; risk flat
T=9–14 Seed error “4 weeks” $\rightarrow$ “40 days” Slight rise in data-driven signal
T=15–22 Propagation “3 $\times$ 40 = 120” Reasoning-driven share begins to rise
T=23–40 Amplification Final answer: \$108 Reasoning-driven dominates (snowballing)
Table 8: Evolution of hallucination in GSM8K arithmetic reasoning.
Case Study 2 — Long-Document Summarization: Misalignment $\rightarrow$ Overreach $\rightarrow$ Fabrication.
Task: Summarize a 5,000-token policy document
Ground truth: Security audit exception applies only to specific log types.
Length (T) Model Behavior HalluGuard Response
T=1–20 Accurate extraction Correct recovery of retention rules Low risk; strong alignment
T=21–40 Misbinding Incorrect merge of distant sections Data-driven signal increases
T=41–95 Drift Overgeneralized suspension claim Reasoning-driven share rises
T=96–170 Fabrication New false rule introduced Reasoning-driven dominates
Table 9: Evolution of hallucination in long-document summarization.
Appendix D Usage of LLM
Large language models (LLMs) were employed in a limited and transparent manner during the preparation of this manuscript. Specifically, LLMs were used to assist with linguistic refinement, style adjustments, and minor text editing to improve clarity and readability. They were not involved in formulating the research questions, designing the theoretical framework, conducting experiments, or interpreting results. All scientific contributions—including conceptual development, methodology, analyses, and conclusions—are the sole responsibility of the authors.